### v2

# Integrable Cosmological Models with an Arbitrary Number of Scalar Fields

Abstract
We consider cosmological models with an arbitrary number of scalar fields nonminimally coupled to gravity and construct new integrable cosmological models. In the constructed models, the Ricci scalar is an integral of motion irrespectively of the type of metric. The general solutions of evolution equations in the spatially flat FLRW metric have been found for models with the quartic potentials.
PACS: 98.80.Cq; 04.50.Kd; 04.20.Jb
1 Introduction
Recent observations show with high accuracy that the Universe is isotropic and homogeneous at large scales. For this reason, cosmological models with scalar fields play a central role in describing the global evolution of the Universe. There are a lot of models with one or a few scalar fields that describe different epochs of the Universe evolution.
The Universe is spatially flat on large scales, so, the spatially flat Friedmann–Lemaître–Robertson–Walker (FLRW) metric with
$ds^{2}={}-N^{2}(\tau)d\tau^{2}+a^{2}(\tau)\left(dx^{2}+dy^{2}+dz^{2}\right),$ (1)
where $N(\tau)$ is the lapse function and $a(\tau)$ is the scale function of the parametric time $\tau$, can be used to describe the global Universe evolution at the background level.
In FLRW metric, the Einstein’s equations reduce to a system of ordinary differential equations, but this system is integrable only for a few specific forms of the potential. The problem of constructing integrable cosmological models is actively investigated, but even taking into account single-field cosmological models, the number of integrable models is not large [1, 2, 3, 4, 5, 6, 7, 8, 9]. In the case of a few scalar fields, even construction of cosmological models with exact particular solutions is not a trivial task [10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
There are a few methods for integrating equations of evolution. The integrability of many cosmological models has been proven by solving evolution equations with suitable parametric time [4], but it is not clear how to find this parametric time. A method for constructing integrable models with non-minimally coupled scalar fields by using the relation between the Jordan and Einstein frames has been investigated for single-field models in Refs. [5, 7, 20]. The integrability of cosmological models with non-minimal coupling can be more apparent than the integrability of the corresponding models in the Einstein frame [7].
The goal of this paper is to construct new integrable cosmological models with a few scalar fields that have the Ricci scalar as an integral of motion. One field cosmological model with such a property has been constructed in Ref. [6]. We construct models with an arbitrary number of non-minimally coupled scalar fields. For constructed models, we find the general solutions of evolution equations in the spatially flat FLRW metric (1).
2 Model with $N$ scalar fields and nonminimal coupling
Let us consider the model, described by the following action
$S=\int d^{4}x\sqrt{-g}\left(U\left(\phi^{A}\right)R-\frac{1}{2}g^{\mu\nu}\delta_{AB}\partial_{\mu}\phi^{A}\partial_{\nu}\phi^{B}-V\left(\phi^{A}\right)\right),$ (2)
where capital Latin indices enumerate fields, $A=1,2,\dots N$, and are lowered and raised with the Kronecker symbol $\delta_{AB}$. The functions $U(\phi^{A})$ and $V(\phi^{A})$ are twice-differentiable functions, $R$ is the Ricci scalar.
The corresponding Einstein’s equations are [8]
$U\left(R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\right)=\nabla_{\mu}\nabla_{\nu}\,U-g_{\mu\nu}\Box\,U+\frac{1}{2}T_{\mu\nu},$ (3)
where
$T_{\mu\nu}=\delta_{AB}\partial_{\mu}\phi^{A}\partial_{\nu}\phi^{B}-g_{\mu\nu}\left(\frac{1}{2}g^{\alpha\beta}\delta_{AB}\partial_{\alpha}\phi^{A}\partial_{\beta}\phi^{B}+V\right).$ (4)
The field equations have the following form
$\Box\phi^{A}=V_{,\phi^{A}}-RU_{,\phi^{A}},$ (5)
where $F_{,\phi^{A}}=\partial F/\partial\phi^{A}$ for any function $F$.
From Eq. (3), we get
$3\Box U-UR=\frac{1}{2}g^{\mu\nu}T_{\mu\nu}={}-\frac{1}{2}g^{\mu\nu}\delta_{AB}\partial_{\mu}\phi^{A}\partial_{\nu}\phi^{B}-2V.$ (6)
Writing $\Box U$ explicitly,
$\Box{U}=U_{,\phi^{A}}\Box\phi^{A}+U_{,\phi^{A}\phi^{B}}g^{\alpha\beta}\partial_{\alpha}\phi^{A}\partial_{\beta}\phi^{B},$ (7)
we obtain from Eq. (6)
$\left(3U_{,\phi^{A}}^{2}+U\right)R-\left[3U_{,\phi^{A}\phi^{B}}+\frac{\delta_{AB}}{2}\right]g^{\alpha\beta}\partial_{\alpha}\phi^{A}\partial_{\beta}\phi^{B}=3U_{,\phi^{A}}V_{,\phi^{A}}+2V.$ (8)
It is easy to see that for the function
$U=U_{0}-\frac{\delta_{AB}}{12}\left(\phi^{A}-\phi_{0}^{A}\right)\left(\phi^{B}-\phi_{0}^{B}\right),$ (9)
where $\phi_{0}^{A}$ is an arbitrary constant vector, Eq. (8) takes the following form:
$\frac{1}{2}\left(\phi^{A}-\phi_{0}^{A}\right)V_{,\phi^{A}}-2V+U_{0}R=0.$ (10)
If $R$ is a constant, $R=R_{0}$, then Eq. (10) is a linear first-order partial differential equation with $V(\phi^{A})$ as an unknown function and the components of $\phi^{A}$ as independent variables. Equation (10) has the general solution
$V=\frac{R_{0}U_{0}}{2}+\left(\phi^{1}-\phi_{0}^{1}\right)^{4}f\left(\frac{\phi^{2}-\phi_{0}^{2}}{\phi^{1}-\phi_{0}^{1}},\frac{\phi^{3}-\phi_{0}^{3}}{\phi^{1}-\phi_{0}^{1}},\dots,\frac{\phi^{N}-\phi_{0}^{N}}{\phi^{1}-\phi_{0}^{1}}\right),$ (11)
where $f$ is an arbitrary differentiable function.
So, we find that the Ricci scalar $R$ is the integral of motion for models with $N$ scalar fields and functions $U$ and $V$ given by Eqs. (9) and (11), respectively. Note that we do not specify the form of the metric. The vector $\phi_{0}^{A}$ can be put to zero without loss of generality.
We have considered the case of $N$ scalar fields with standard kinetic terms. To generalize this result to the case with a phantom scalar field $\phi^{A}$, it is enough to assume that the scalar field $\phi^{A}$ has the standard kinetic part, but is a purely imaginary function. Note that the function $U$ remains real in this case, and the potential $V$ is real for any real-valued function $f$.
3 General solutions in the FLRW metric
In the FLRW metric (1), Eqs. (3) and (5) take the following form:
$6Uh^{2}{}+6h\dot{U}=\frac{1}{2}\delta_{AB}\dot{\phi}^{A}\dot{\phi}^{B}+N^{2}V,$ (12)
${}-U\left(2\dot{h}+3h^{2}\right)-2Uh\frac{\dot{N}}{N}=\ddot{U}+2h\dot{U}-\dot{U}\frac{\dot{N}}{N}+\frac{1}{2}\left(\frac{1}{2}\delta_{AB}\dot{\phi}^{A}\dot{\phi}^{B}-N^{2}V\right),$ (13)
$\ddot{\phi}^{A}+\left(3h-\frac{\dot{N}}{N}\right)\dot{\phi}^{A}+V_{,\phi^{A}}N^{2}-6\left(\dot{h}+2h^{2}-h\frac{\dot{N}}{N}\right)U_{,\phi^{A}}=0,$ (14)
where $h(\tau)\equiv\dot{a}/a$, dots mean derivatives with respect to the parametric time $\tau$, and we assume that all fields are homogeneous.
For the function $U$ given by (9), the field equations (14) take the form
$\ddot{\phi}^{A}+\left(3h-\frac{\dot{N}}{N}\right)\dot{\phi}^{A}+V_{,\,\phi^{A}}N^{2}+\left(\dot{h}+2h^{2}-h\frac{\dot{N}}{N}\right)\phi^{A}=0$ (15)
To solve Eq. (15), we use the conformal time $\eta$, defined as $dt=a(\eta)d\eta$, hence, $N(\eta)=a(\eta)$. In terms of functions $y^{A}(\eta)\equiv a(\eta)\ \phi^{A}(\eta)$, the field equations (15) are
$\frac{d^{2}y^{A}}{d\eta^{2}}+a^{3}V_{,\phi^{A}}\left(\frac{y^{1}}{a},\dots,\frac{y^{N}}{a}\right)=0.$ (16)
Moreover, if we take $V$ as in (11), we get
$\frac{d^{2}y^{A}}{d\eta^{2}}+\frac{\partial}{\partial y^{A}}V\left(y^{1},\dots,y^{N}\right)=0,$ (17)
where
$V\left(y^{1},\dots,y^{N}\right)=\frac{1}{2}R_{0}U_{0}+\left(y^{1}\right)^{4}f\left(\frac{y^{2}}{y^{1}},\frac{y^{3}}{y^{1}},\dots,\frac{y^{N}}{y^{1}}\right).$
The system (17) has the first integral
$\frac{1}{2}\delta_{AB}\frac{dy^{A}}{d\eta}\frac{dy^{B}}{d\eta}+V\left(y^{1},\dots,y^{N}\right)=\frac{1}{2}R_{0}U_{0}+E,$ (18)
where $E$ is an integration constant.
For the previously chosen functions $U$ and $V$, the constraint equation (12) in the conformal time takes the following form
$\left(\frac{h}{a}\right)^{2}=\frac{R_{0}}{12}+\frac{E}{6U_{0}a^{4}}.$ (19)
Let’s consider a particular kind of the potential $V$:
$V(\phi^{1},\dots,\phi^{N})=\frac{R_{0}U_{0}}{2}+\sum_{A=1}^{N}c_{A}\left(\phi^{A}\right)^{4},$ (20)
where $c_{A}$ are constants.
With such a potential, system (17) is completely separable, and we have $N$ independent integrals of motion:
$\frac{1}{2}\left(\frac{dy^{A}}{d\eta}\right)^{2}+c_{A}\left(y^{A}\right)^{4}=C^{A},$ (21)
where $C^{A}$ are integration constants, and the Einstein summation convention is not applied to $A$. We get that $E=\sum\limits_{A=1}^{N}C^{A}$, and
$H^{2}(\tau)=\left(\frac{h}{a}\right)^{2}=\frac{R_{0}}{12}+\frac{1}{6U_{0}a^{4}}\sum_{n=1}^{N}C^{A}.$ (22)
For the lapse function $N=1$, the condition $R=R_{0}$ gives the first order differential equation in the Hubble parameter $H(t)$,
$\dot{H}+2H^{2}=\frac{R_{0}}{6},$ (23)
that has the following general solution at $R_{0}>0$:
$H(t)=\frac{\sqrt{3R_{0}}\left(\mathrm{e}^{2\sqrt{3R_{0}}t/3}+B\right)}{6\left(\mathrm{e}^{2\,\sqrt{3R_{0}}t/3}-B\right)}\,,$ (24)
where $B$ is a constant.
There are several different solutions for $H$, $a$ and $\eta$, depending on the sign of $E$ and on whether or not $R_{0}$ is equal to zero. Since the fields $y_{n}$ are most simply written as functions of $\eta$, it is helpful to write out $H$ and $a$ as functions of $\eta$ too, as well as to explicitly write the domain of $\eta$.
For $R_{0}>0$, we obtain
  1. 1.
    if $E>0$, then
    $\begin{split}&H(t)=\sqrt{\frac{R_{0}}{12}}\coth\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right),\\ &a(t)=a_{0}\sqrt{\sinh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right)},\end{split}$ (25)
    $\eta(t)=\eta_{0}+\frac{1}{a_{0}}\sqrt{\frac{3}{R_{0}}}F\left(\arccos\left(\frac{a_{0}^{2}-a^{2}}{a_{0}^{2}+a^{2}}\right)\left|\frac{1}{2}\right.\right);$
    $H(\eta)=\sqrt{\frac{R_{0}}{3}}\frac{\mathrm{dn}\left(\left.a_{0}\sqrt{\frac{R_{0}}{3}}\left(\eta-\eta_{0}\right)\,\right|\frac{1}{2}\right)}{1-\mathrm{cn}\left(\left.a_{0}\sqrt{\frac{R_{0}}{3}}\left(\eta-\eta_{0}\right)\,\right|\frac{1}{2}\right)},$ (26)
    $a(\eta)=\frac{a_{0}\,\mathrm{sn}\left(\left.a_{0}\sqrt{\frac{R_{0}}{3}}\left(\eta-\eta_{0}\right)\,\right|\frac{1}{2}\right)}{1+\mathrm{cn}\left(\left.a_{0}\sqrt{\frac{R_{0}}{3}}\left(\eta-\eta_{0}\right)\,\right|\frac{1}{2}\right)},\quad a_{0}=\left(\frac{2E}{U_{0}R_{0}}\right)^{\frac{1}{4}},$ (27)
    $a_{0}\sqrt{\frac{R_{0}}{12}}\left(\eta-\eta_{0}\right)\in\left(0,K\left(\frac{1}{2}\right)\right);$
  2. 2.
    if $E<0$, then
    $\begin{split}&H(t)=\sqrt{\frac{R_{0}}{12}}\tanh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right),\\ &a(t)=a_{0}\sqrt{\cosh\left(\sqrt{\frac{R_{0}}{3}}(t-t_{0})\right)},\end{split}$ (28)
    $\eta(t)=\eta_{0}+\frac{1}{a_{0}}\sqrt{\frac{6}{R_{0}}}F\left(\arcsin\left(\sqrt{\frac{a^{2}-a_{0}^{2}}{a^{2}}}\right)\left|\frac{1}{2}\right.\right);$
    $a(\eta)=\frac{a_{0}}{\mathrm{cn}\left(\left.a_{0}\sqrt{\frac{R_{0}}{6}}\left(\eta-\eta_{0}\right)\,\right|\frac{1}{2}\right)},\quad a_{0}=\left(\frac{-2E}{U_{0}R_{0}}\right)^{\frac{1}{4}},$ (29)
    $H(\eta)=\sqrt{\frac{R_{0}}{6}}\mathrm{sn}\left(\left.a_{0}\sqrt{\frac{R_{0}}{6}}\left(\eta-\eta_{0}\right)\,\right|\frac{1}{2}\right)\mathrm{dn}\left(\left.a_{0}\sqrt{\frac{R_{0}}{6}}\left(\eta-\eta_{0}\right)\,\right|\frac{1}{2}\right),$ (30)
    $a_{0}\sqrt{\frac{R_{0}}{6}}\left(\eta-\eta_{0}\right)\in\left(-K\left(\frac{1}{2}\right),K\left(\frac{1}{2}\right)\right);$
  3. 3.
    if $E=0$, then
    $H(t)=\pm\sqrt{\frac{R_{0}}{12}},\qquad a(t)=a_{0}e^{\pm\sqrt{R_{0}/12}(t-t_{0})},$ (31)
    $\eta(t)=\eta_{0}\pm\frac{1}{a_{0}}\sqrt{\frac{12}{R_{0}}}\left(1-e^{\mp\sqrt{R_{0}/12}(t-t_{0})}\right)=\eta_{0}\pm\frac{1}{a_{0}}\sqrt{\frac{12}{R_{0}}}\left(1-\frac{a_{0}}{a}\right);$
    $H(\eta)=\pm\sqrt{\frac{R_{0}}{12}},\qquad a(\eta)=\frac{a_{0}}{1\mp a_{0}\sqrt{R_{0}/12}\left(\eta-\eta_{0}\right)},$ (32)
    ±$\pm a_{0}\sqrt{\frac{R_{0}}{12}}\left(\eta-\eta_{0}\right)\in\left(-\infty,1\right);$
  4. 4.
    if $R_{0}=0$ and $E>0$, then
    $H(t)=\frac{1}{2(t-t_{0})},\quad a(t)=a_{0}\sqrt{t-t_{0}},$ (33)
    $\eta(t)=\eta_{0}+\frac{2}{a_{0}}\sqrt{t-t_{0}}=\eta_{0}+\frac{2}{a_{0}^{2}}a.$
    $H(\eta)=\frac{2}{a_{0}^{2}(\eta-\eta_{0})^{2}},\quad a(\eta)=\frac{1}{2}a_{0}^{2}\left(\eta-\eta_{0}\right),\quad a_{0}=\left(\frac{2E}{3U_{0}}\right)^{\frac{1}{4}},$ (34)
    $\eta-\eta_{0}\in\left(0,+\infty\right).$
Here $F$ is the incomplete elliptic integral of the first kind, defined as
$F\left(\left.\varphi\,\right|m\right)=\int\limits_{0}^{\varphi}\frac{d\theta}{\sqrt{1-m\sin^{2}\theta}},$ (35)
$K(m)\equiv F(\pi/2\left|m\right.)$ is the complete elliptic integral of the first kind, and also the Jacobi elliptic functions $\mathrm{sn}$, $\mathrm{cn}$ and $\mathrm{dn}$ were introduced. By definition, if $u=F\left(\left.\varphi\,\right|m\right)$, then
$\begin{array}[]{ll}\mathrm{am}\left(\left.u\,\right|m\right)=\varphi,&\qquad\mathrm{dn}\left(\left.u\,\right|m\right)=\frac{d}{du}\mathrm{am}\left(\left.u\,\right|m\right),\\ \mathrm{sn}\left(\left.u\,\right|m\right)=\sin\left(\mathrm{am}\left(\left.u\,\right|m\right)\right),&\qquad\mathrm{cn}\left(\left.u\,\right|m\right)=\cos\left(\mathrm{am}\left(\left.u\,\right|m\right)\right).\\ \end{array}$ (36)
The form of solutions of (21) depends on the signs of $c_{A}$ and $C^{A}$:
  1. 1.
    $C^{A}>0,\quad c_{A}>0$:
    $y^{A}(\eta)=\left(\frac{C^{A}}{c_{A}}\right)^{\frac{1}{4}}\mathrm{cn}\left.\left(\pm 2\left(C^{A}c_{A}\right)^{\frac{1}{4}}(\eta-\eta_{i})+\mathrm{cn}^{-1}\left(u_{A}\left|\frac{1}{2}\right.\right)\right|\frac{1}{2}\right);$
  2. 2.
    $C^{A}<0,\quad c_{A}<0$:
    $y^{A}(\eta)=\frac{\left(\frac{C^{A}}{c_{A}}\right)^{\frac{1}{4}}}{\mathrm{cn}\left.\left(\pm 2\left(C^{A}c_{A}\right)^{\frac{1}{4}}(\eta-\eta_{i})+\mathrm{cn}^{-1}\left.\left(\frac{1}{u_{A}}\right|\frac{1}{2}\right)\right|\frac{1}{2}\right)};$
  3. 3.
    $C^{A}>0,\quad c_{A}<0$:
    $y^{A}(\eta)=\left|\frac{C^{A}}{c_{A}}\right|^{\frac{1}{4}}\frac{\mathrm{sn}\left(\pm\left.2\sqrt{2}\left|C^{A}c_{A}\right|^{\frac{1}{4}}(\eta-\eta_{i})+\mathrm{sn}^{-1}\left(\left.\frac{2u_{A}}{1+u_{A}^{2}}\right|\frac{1}{2}\right)\right|\frac{1}{2}\right)}{1+\mathrm{cn}\left(\pm\left.2\sqrt{2}\left|C^{A}c_{A}\right|^{\frac{1}{4}}(\eta-\eta_{i})+\mathrm{cn}^{-1}\left(\left.\frac{1-u_{A}^{2}}{1+u_{A}^{2}}\right|\frac{1}{2}\right)\right|\frac{1}{2}\right)}.$
    Constants $u_{A}=|c_{A}/C_{A}|^{1/4}y^{A}(\eta_{i})$ and the choice of sign in “$\pm$” are determined by initial conditions set at $\eta=\eta_{i}$. Note that no summation over $A$ is assumed.
  4. 4.
    For $c_{A}=0$, $C^{A}$ has to be non-negative, and we have
    $y^{A}(\eta)=y^{A}(\eta_{i})+\frac{dy^{A}}{d\eta}(\eta_{i})(\eta-\eta_{i}).$
  5. 5.
    There are no solutions when $C^{A}<0$ and $c_{A}\geqslant 0$.
Note that the bounce solution (28) exists only if at least one of $c_{A}<0$ and, therefore, the potential $V$ is not bounded from below.
4 Conclusions
In this paper, new integrable models with $N$ scalar fields non-minimally coupled to gravity have been constructed. We have shown that the Ricci scalar is an integral of motion if the functions $U$ and $V$ are given by Eqs. (9) and (11). In the spatially flat FLRW metric, the Hubble parameter can be presented in terms of hyperbolic functions of the cosmic time. The behaviour of the scalar fields have been described in terms of the elliptic functions of the conformal time.
The corresponding one field model proposed in Ref. [6] can be transformed to the integrable model with a scalar field minimally coupled to gravity [7]. In general case, it is not possible to transform models with a few scalar fields non-minimally coupled to gravity into models with minimally coupled scalar fields with standard kinetic terms [21]. After the metric transformation, one obtains models with a non-standard kinetic part, the so-called chiral cosmological models [22, 23, 15].
The simplest case of an integrable chiral cosmological model is the two-field model with a constant potential. The absence of the scalar field dependence of the potential allows one to obtain a description of the behaviour of the Hubble parameter in the form of hyperbolic functions [24].
New integrable models with minimally coupled scalar fields can be obtained by conformal metric transformation of integrable models with non-minimally coupled scalar fields [25]. Another possible way to generalize the obtained results is to get the general solutions of integrable FLRW models in open and closed FLRW metrics [7, 26] or in the Bianchi I metric [8, 20, 27].
Funding
V.R.I. is supported by the “BASIS” Foundation grant No. 22-2-2-6-1.